OpenStudy (eriyalovesu):

HELP WILL MEDAL! Solve the following system: y = x + 3 3x + y = 19 (7, 4) (−4, 7) (4, 7) (4, −7)

OpenStudy (eriyalovesu):

@agent0smith

OpenStudy (mayankdevnani):

plug the 1st equation in second equation

OpenStudy (eriyalovesu):

for x?

OpenStudy (mayankdevnani):

for y

OpenStudy (eriyalovesu):

(3x + y = 19)= x + 3

OpenStudy (mayankdevnani):

nope not like that let me show you how

OpenStudy (mayankdevnani):

3x+y=19 and given that y=x+3 so,we get 3x+(x+3)=19

OpenStudy (mayankdevnani):

did you get it?

OpenStudy (eriyalovesu):

yes no combine like terms

OpenStudy (eriyalovesu):

now

OpenStudy (mayankdevnani):

yea

OpenStudy (eriyalovesu):

no distributive property?

OpenStudy (mayankdevnani):

nope cuz they are adding

OpenStudy (eriyalovesu):

4x+3=19

OpenStudy (mayankdevnani):

good

OpenStudy (mayankdevnani):

now subtract both sides by 3

OpenStudy (eriyalovesu):

4x=16

OpenStudy (mayankdevnani):

good

OpenStudy (mayankdevnani):

now divide both sides by 4

OpenStudy (eriyalovesu):

x=4

OpenStudy (mayankdevnani):

good

OpenStudy (mayankdevnani):

now plug x=4 in the equation y=x+3

OpenStudy (mayankdevnani):

find y from there

OpenStudy (eriyalovesu):

y=7

OpenStudy (mayankdevnani):

awesome

OpenStudy (mayankdevnani):

there you go ;)

OpenStudy (mayankdevnani):

option C looks correct

OpenStudy (eriyalovesu):

Thank you!

OpenStudy (mayankdevnani):

np

OpenStudy (eriyalovesu):

@mayankdevnani can you help with one more?

OpenStudy (mayankdevnani):

sure @eriyalovesyou

OpenStudy (mayankdevnani):

@eriyalovesu